OpenStudy (anonymous):

evaluate the integral 1/(x^(2)sqrt(4-x^(2))

OpenStudy (anonymous):

Is this what you intend?\[\int \dfrac{dx}{x^2 \sqrt{4-x^2}}\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Try the substitution:\[x=2\sin u\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok :) let me try it from here

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok so i got it simplified down to: \[\int\limits_{?}^{?}1\div \sin ^{2}xcosx du \]

OpenStudy (anonymous):

do i do a u subsitution?

OpenStudy (anonymous):

You forgot that\[dx = 2\cos u \text{ } du\]